Two sounds of equal intensity I generates beats. The max. intensity of sound produced in beats will be:


A)I


B)2I


C)4I


D)I/2


 


Thanks!!


Grade:11

3 Answers

Prakhar Srivastava
11 years ago

Dear Prajwal kr

intensity max= ((I1)1/2+(I2)1/2)2  AS I1=I2

therfore by the realtion we get (I+I)2=4I

C)IS CORRECT..!!!

Thnks..!!

Roshan Mohanty
11 years ago

I'' =  root(I1^2) + root(I2^2) +2root(I1)root(I2)cos @
   =root(I^2) + root(I^2)
   =2I 

Please approve :) 

Shivam Dimri
11 years ago

One very important fact

two non coherent sources will never give a clear interference pattern

hence you only talk about the average of them

maxium = minimum = I + I = 2I

always...